OpenStudy (anonymous):

solve for x ln(x^2)=4

jimthompson5910 (jim_thompson5910):

\[\large \ln(x^2)=4\] \[\large x^2=e^4\] \[\large x=\pm\sqrt{e^4}\] \[\large x=\pm e^2\] \[\large x=e^2 \ \ \text{or} \ \ x = -e^2\]